15-16 October 2008

Observers: , DThompson, DMiller, ARakich, JHill (Tucson), LUCIFER folks (N.Ageorges,W.Seifert, et al)
Software Support: TSargent in Tucson
Telescope Operators: A.Ceranski

Summary

Dave et al started up the telescope with a few problems, but by 7:30 the telescope was running well. The rest of the night we only had GCS freeze (waiting on image) twice. Otherwise, no problems.

We took field aberration data at a high elevation (72->62) and at low elevation (25->34) and neither showed much binodal astigmatism. It thus looks like there is no dependence of binodal astigmatism on elevation. A collimation lookup table for the secondary with constant values looks like a good first approximation. Thus, if a new collimation and pointing model need to be measured, the position of the secondary for all elevations is good.

Lucifer was still seeing some astigmatism in there images. I first look seems the the astigmatism is about the same at all (25->80) elevations. We took some through focus image so Andrew and get quantitative amount of astigmatism.

Lucifer took some sky flats with the morning sky.

Details

HBS and other problems at the beginning of the night.

UT01:57 Telescope pointing corrected. 12.5C outside temp. M1=9.2C

BS9185 - acquired OK. EL=50. WFS sent 1800nm on first correction. Guider is seeing 0.6 arcsec, LUCIFER is seeing 0.3 arcsec. WFS converged in a few cycles from a severely comatic image.

BS9182 - acquired OK, EL=52. Initial collimation good. After collimation the guider sees 0.55", the WFS is sending ~200nm rms.

BS9177 - Acquisition failed (image 160), EL=62. The star The guider is driving us off the source. Checked pointing, IE was off by 12 arcsec...why? Source was on the CCD.

BS9177 - acquired OK, EL=58. Source centered on acquisition image. Guider is seeing 0.6 arcsec. Walter reports LUCIFER sees some residual coma on LUCIFER.

MST from now on! DLM

8:00 Lucifer preset to BS9188, GUIDE and Collimate

8:15 Doug send preset to BS9188, change to IRTC so pointing offsets from earlier in the night. No star on Acquisition image. Re-authorize Lucifer and Walter reset the pointing model offset.

Field Measurement

8:45 Doug send preset to BS9188 as Lucifer. Run ActOpt to collimate wfsc 5-11

9:02 gstar = 24 PA=189 wfsc 12-14

9:09 on-axis PA=189 wfsc 15-16

9:13 gstar = 32 PA=61 wfsc 17-19

9:20 on-axis PA=61 no guide star found...found a patch of noise on acquire image and then couldn't guide

9:30 edit

9:35 on-axis PA=61 wfsc 20-22

9:38 gstar=10 PA=315 wfsc 23-25

9:42 on-axis PA=315 wfsc 26-29

9:50 !!John noticed that we were sending focus and coma to secondary. John clear Act Opt corrections on Secondary

9:51 collimate on axis (high coma) wfsc 30-37

10:05 gstar=24 PA=189 star not on acquisition

10:07 Walter and Nancy re-center field on Lucifer

10:15 IE=? CA=?

10:40 Lucifer finish with alignment

10:41 on-axis PA=315 wfsc 38-40

10:44 gstar 10 PA=315 error

22:50 GCS froze with "waiting for image acquisition" printed twice. Aaron stop and start GCS

10:52 on-axis PA=315 wfsc 41-42

10:54 gstar=10 PA=315 wfsc 43-45 (grid problem?) 46 ok, 47 bad, 48 ok, 49 ok

11:03 on-axis PA=315 Couldn't find guide star

11:10 Lucifer re-center star field

11:15 on-axis PA=315 wfsc 50-51

11:20 gstar=14 PA=230 wfsc 52-54

11:23 on-axis PA=230 wfsc 55-57

11:27 gstar=33 PA=114 wfsc 58-60

11:31 on-axis PA=114 wfsc 61-63

11:35 gstar=30 PA=45 wfsc 64-66

11:38 on-axis PA=45 wfsc 67-69

11:43 gstar=10 PA=315 wfsc 70-72

11:46 on-axis PA=315 wfsc 73

11:51 gstar=23 PA=199 wfsc 74-76

11:56 on-axis PA=199 wfsc 77-79

12:02 gstar=32 PA=61 wfsc 80-82

12:06 on-axis PA=61 wfsc 83-84

12:10 gstar=7 PA=276 wfsc 85-87

12:15 on-axis PA=276 wfsc 88-89

Lucifer field measurements

12:22 gstar=32 PA=61 collimate wfsc 90-91

12:27 Bumped into Y limit: adjust M1 global Y value to +1 and M2 global y value to -2.766. Guiding continued

12:29 collimate wfsc 92-99

12:38 Nancy take images with Lucifer on

focus name (uci_20081015_000*.fits)
0 6
2000 7
1000 8
0 9
-1000 10
-2000 11
0 12

wfsc 100 not applied

12:52 gstar=22 PA=338 collimate wfsc 101-104

12:38 Nancy take images with Lucifer on

focus name (uci_20081015_000*.fits)
0 13
2000 14
1000 15
0 16
-1000 17
-2000 18
0 19

1:08 gstar=22 PA=338 collimate 105-108

1:14 on-axis PA=338 wfsc 109-110

1:19 gstar=32 PA=61 collimate 111-113

1:23 on-axis PA=61 wfsc 114-115

1:30 DIMM fellows need to reset a box, so Aaron tips the telescope over to horizon.

1:40 Telescope is back to previous position

1:40 preset on-axis BS9188, ACTIVE collimate 46-52

1:50 preset gstar=33 , ACTIVE failed

1:55 Aaron stop and start GCS "waiting for image to arrive"

1:56 preset gstar=33, ACTIVE, 53-59

2:02 Lpoint EL30on_axis GUIDE

Field Aberrations Measurement at low elevation

2:45 pointto BS9134 "ACTIVE" collimate throught 73

3:03 on-axis PA=0 collimate wfsc 116-117

3:05 gstar=37 PA=83 wfsc 118-120

3:12 on-axis PA=83 wfsc 121-122

3:18 gstar=28 PA=349 wfsc 123 ok 124-125 bad 126-127 ok

3:22 on-axis PA=349 wfsc 128-130

3:26 gstar=12 PA=261 wfsc 131-133

3:32 on-axis PA=261 wfsc 134-136

3:37 gstar=27 PA=198 wfsc 137-139

3:43 on-axis PA=198 wfsc 140-141

3:46 gstar=40 PA=108 wfsc 142-144

3:48 on-axis PA=108 wfsc 145-146

3:53 gstar=31 PA=15 wfsc 147-149

3:57 on-axis PA=15 wfsc 150-151

4:00 gstar=11 PA=295 wfsc 152-154

4:04 on-axis PA=295 wfsc 155-157

4:08 gstar=17 PA=212 wfsc 158-160

4:15 on-axis PA=211 wfsc 161-162

4:25 gstar=40 PA=180 wfsc 163-165

Field aberrations measured by Lucifer

5:02 pointto BS9105 GUIDE (elevation 30)

5:02 on-axis collimate wfsc 170

5:04 gstar=3 PA=262 collimate wfsc 171-176

Focus Name (luci_20081015_000*.fits)
0 45
2000 46
1000 47
0 48
-1000 49
-2000 50
0 51

5:24 gstar=6 PA=337 collimate wfsc 177-179

Focus Name (luci_20081015_000*.fits)
0 52
2000 53
1000 54
0 55
-1000 56
-2000 57
0 58

5:36 BS9134 gstar=40 PA=108 wfsc 180

5:39 M1 bump against the y limit so Andrew set m2 y global offset to -1.766, m1 y global offset set to +2

5:41 gstar=40 wfsc 181-189

0 59
2000 60
-2000 63

5:56 gstar=40 PA=153 wfsc 190-191

0 59
2000 60
-2000 63

-- DavidThompson - 09 Oct 2008
Topic revision: r7 - 19 Oct 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback